# Leveraging Semantic Attribute Binding for Free-Lunch Color Control in Diffusion Models

### v1

@@ 1 Introduction
Recent advances in text-to-image (T2I) diffusion models [40, 35, 39] have revolutionized image generation, offering unprecedented control over various attributes including composition [26, 49], style [42, 45, 50], subject matter [10, 38], and visual aesthetics [13, 12]. These models can generate complex scenes with specific objects, artistic styles, and spatial arrangements based on textual descriptions. Among these controllable attributes, color stands as particularly crucial, as it fundamentally shapes viewer perception, conveys emotional tone, and often serves as a critical design specification in professional contexts from brand identity to product development [41]. Yet despite remarkable progress in generative capabilities, T2I models exhibit a significant limitation in precise color control. When users specify colors through linguistic descriptors like “red” or “blue”, these terms encompass broad ranges of potential shades, making it challenging to achieve exact color matching in generated content—a capability essential for practical design applications where color fidelity is non-negotiable.
@@
T2I personalization methods, such as DreamBooth [38] or CustomDiffusion [21], can be used to address this limitation and learn precise color prompts. However, these methods were shown to obtain unsatisfactory color generations [4], often failing to correctly disentangle shape from color information. Methods such as ColorPeel [4] enable precise color control by learning specific color prompts through additional training. However, these approaches require separate optimization processes for each individual color, resulting in significant computational overhead and limiting practicality. The need for model fine-tuning creates a fundamental barrier to flexibly specifying arbitrary colors during inference time, constraining creative applications and design tasks that demand precise color control. Furthermore, the learning-based strategy of ColorPeel becomes harder to implement in advanced T2I diffusion models, where multiple text encoders are commonly employed to process textual inputs. Models such as SDXL [33], SD3 [9], and Flux [22] introduce additional complexity, making training-based methods increasingly impractical and often ineffective for precise color control.
@@
Recent advancements in image conditioning for diffusion models, particularly through mechanisms like IP-Adapter [51, 45, 50], have enabled new forms of control by conditioning generation on reference images. While these adapters were primarily designed for style transfer or subject-driven generation, they encode complex visual information in ways that have not been fully explored. Through systematic investigation, we discovered a previously unrecognized property: semantic attribute binding which refers to the implicit bindings between visual attributes and their semantic representations within multi-modal space. Specifically, we found that these bindings create a mapping between RGB values in reference images and linguistic color descriptors in text prompts—a connection that existing methods, including IP-Adapter itself, do not explicitly exploit. Our insight enables precise RGB-level color manipulation without any model fine-tuning, revealing capabilities latent within IP-Adapter that haven’t been previously identified or utilized for targeted control.
@@ 2.1 Image-Conditioned T2I Diffusion Models
Conditional image-based text-to-image (T2I) models integrate auxiliary image inputs with textual prompts, allowing for more refined control over spatial layout [24, 7], object content [6, 23], and stylistic attributes [42, 14]. For pixel-level structural guidance, ControlNet [52] employs a duplicated U-Net [36] encoder to process various input maps, while UniControlNet [54] extends this with dual adapters handling both local (pixel-level) and global (CLIP image embedding) controls, enabling simultaneous spatial and content-aware guidance.
@@
As another technique pipeline, IP-Adapter [51] leverages the Vision Transformer (ViT) encoder from CLIP [34] to extract real image features, integrating them into the U-Net backbone via cross-attention layers for more coherent conditioning. Building on this, IPAdapter-Instruct [37] introduces an instructional prompt mechanism, enhancing interpretability and reducing the need for separate model training for different conditioning inputs, addressing a key limitation of ControlNet. However, the role of the additional cross-attention mechanism in IP-Adapter remains underexplored. In this paper, we present the first findings on its properties, revealing that IP-Adapter implicitly establishes a binding between visual attributes and their corresponding semantic representations.
2.2 Color Control in T2I diffusion models
@@
With the rapid advancements in T2I generation, various text-guided image editing approaches [15, 29, 30, 53, 5, 25, 47] have been developed to enable controllable modifications. For instance, methods like Imagic [18] and P2P [16] leverage Stable Diffusion (SD) models for structure-preserving edits. InstructPix2Pix [3] is an extension of P2P by allowing human-like instructions for image editing. Following that, pix2pix-zero [32] propose noise regularization and cross-attention guidance to retrain the structure of a given image. Another technique stream which can also achieve controllable generation is transfer learning for T2I models, or also referred to T2I model adaptation and personalized generation. It aims at adapting a given model to a new concept by given images from the users and bind the new concept with a unique token. As a result, the adaptation model can generate various renditions for the new concept guided by text prompts. Depending on whether the adaptation method is fine-tuning the T2I model [38, 21], or freezing the T2I backbone [10, 48, 8].
@@
However, all these existing techniques rely heavily on the generative capacity of diffusion models and struggle to achieve fine-grained control over color attributes in image editing and generation tasks. Only a limited number of works [4, 11] have begun addressing the challenge of precise color generation. Rich-Text [11] enhances color fidelity through multi-pass processing with global and local diffusion models, but at the cost of high computational overhead and reduced color accuracy. In contrast, ColorPeel [4] introduces color prompt learning to improve color alignment with user input. However, this method requires extensive training and is restricted to handling only a few color names per training session. In this paper, we argue that the limitations of prior approaches stem from their failure to recognize the implicit attribute binding between the IP-Adapter [51] module and T2I diffusion models [33, 35]. Leveraging this property, we propose a training-free solution for precise color generation, allowing seamless adaptation to any user-specified color inputs, particularly benefiting artistic and creative applications.
@@ 3.1 Preliminaries
T2I Diffusion Models. We built on the SDXL [33] model, consisting of two primary components: an autoencoder and a diffusion model $\epsilon_{\theta}$. The model $\epsilon_{\theta}$ is trained by the loss:
@@
where $\epsilon_{\theta}$ is a UNet, conditioning a latent input $z_{t}$, a timestep $t\sim\text{U}(1,T)$, and a text embedding $\tau_{\xi}(\mathcal{P})$. More specifically, text-guided diffusion models generate an image from the textual condition as $\mathcal{C}_{text}=\tau_{\xi}(\mathcal{P})$, where $\tau_{\xi}$ is the CLIP text encoder [17]  [^1] SDXL uses two text encoders and concatenate the embeddings.. The cross-attention map is derived from $\epsilon_{\theta}(z_{t},t,\mathcal{C}_{text})$. After predicting the noise, diffusion schedulers [27, 28] are used to predict the latent $z_{t-1}$. As an example with the DDIM scheduler [43], the formula is:
@@
IP-Adapter. Building on T2I diffusion models, the IP-Adapter [51] introduces additional controllability by conditioning the T2I model on a conditional image $\mathcal{I}_{ip}$. Practically, this involves leveraging a pretrained T2I diffusion model and incorporating a cross-attention layer to the (projected) image condition following each text-prompt conditioning layer. The conditional image is encoded in the low-dimensional CLIP image embedding space [17] to capture high-level semantic information. By denoting the CLIP image encoder as $\tau_{\phi}$ and IP-Adapter projection as $\mathbf{IP}$, this process is adding a new image condition $\mathcal{C}_{img}=\mathbf{IP}(\tau_{\phi}(\mathcal{I}_{ip}))$ to the T2I model as $z_{t-1}=\mathcal{G}(z_{t},t,\mathcal{C}_{text},\mathcal{C}_{img})$.
@@ 3.2 Semantic attribute binding in IP-Adapters
Figure 2: Illustration of semantic attribute binding. On top, the color-guidance image is provided. a) Given the color name used in the prompt, the generated results will pick a different color from the color-guidance image. b) Changing the colors of the color guidance image, results in similar changes in the generated images. c) The exact color which should be generated for the used color name can also be a synthetic example.
IP-Adapter [51, 37] has demonstrated remarkable effectiveness in transferring visual attributes from reference images to generated content. In this paper, we introduce a key insight: the IP-Adapter implicitly establishes latent-space connections between visual features and their semantic counterparts. As illustrated in Figure 2, this phenomenon, which we call semantic attribute binding, creates a direct correspondence between colors in reference images and their linguistic descriptors in prompts. When using a single reference image containing multiple colors (Figure 2a), the model correctly extracts and applies each color based on the corresponding color name in the prompt (e.g., “green”, “yellow” or “red”). Furthermore, when the colors in the reference image are modified (Figure 2b), the generated results reflect these changes while maintaining the binding to the appropriate color names. This binding even works with synthetic color patches (Figure 2c), allowing for precise RGB-level control. These observations reveal that the adapter inherently associates specific color values with their linguistic descriptors, enabling exact color specification without requiring any additional training or fine-tuning.
Color name attribution.
Semantic attribute binding stems from the architectural design of IP-Adapter, particularly its decoupled cross-attention mechanism. As shown by Ye et al. [51], the IP-Adapter introduces separate cross-attention layers for image features while keeping the original text cross-attention layers intact. Crucially, both cross-attention mechanisms share the same query matrices, while maintaining separate key and value projection matrices for text and image features:
@@
To effectively leverage this semantic attribute binding, we need to determine which color names in natural language correspond to specific RGB values. Following foundational work in color linguistics by Berlin and Kay [2], we recognize that human languages typically use a limited set of basic color terms to describe the entire color spectrum. In English, there are eleven basic color terms, with eight being chromatic (red, green, blue, yellow, brown, orange, pink, and purple).
@@
Interestingly, we observe that the similarity values are not uniform across all color pairs. Some color terms (like “green” and “blue”) show stronger and more distinctive binding patterns, while others exhibit more diffuse relationships. This variation aligns with findings in color linguistics that certain basic color terms have more centralized and consistent representations across languages and visual systems [19].
By quantifying these semantic bindings, we can automatically determine the appropriate color name to use in prompts based on user-specified RGB values, creating a bidirectional mapping between the continuous color space and discrete linguistic color categories. This mapping forms the foundation of our color control approach, allowing us to exploit the implicit color knowledge already encoded in the model’s attention mechanisms.
@@
To address these challenges, our ColorWave approach introduces a selective attention modulation strategy that preserves the generative capabilities of the base model while enabling precise color control for targeted objects.
@@
3.3 Spatial Prior Addition
In our approach ColorWave , we leverage the decoupled cross-attention design established in previous adapters [51, 45] for its minimal interference with pretrained knowledge and compatibility with other methods. We implement a strategy of minimal yet precise intervention in the generative model to achieve effective color binding. First, we inject the user-specified color information at a single optimal point in the network, where it most effectively influences the desired subject, following successful approaches in prior work [45, 46]. Second, we rely on the model’s inherent understanding to determine proper color placement based on its pretrained knowledge.
@@ 3.3 Spatial Prior Addition
$\mathbf{x}=\mathbf{A}+\left\lceil\mathbf{A}_{o}\right\rceil\mathbf{A}^{\prime}.$ (4)
[…]
@@ 4.1 Experimental setup
Dataset. There is no standardized dataset specifically designed for evaluating precise color control in diffusion models. For fair and direct comparison with prior work, we adopt the evaluation framework established by ColorPeel [4], which defines two color generation tasks:
  • Coarse-grained color set: Four basic colors (red, green, blue, and yellow) with predefined RGB values.
  • Fine-grained color set: Eighteen more specialized colors including ”salmon”, ”beige”, ”navy”, and ”indigo”, each with specific RGB values.
For both sets, we use identical prompts to those used in ColorPeel, including objects like “a [color] bowl on the table”, “a [color] teddy bear in Times Square”, and “a [color] sofa in living room”. Each prompt is rendered with 20 different random seeds to account for generation variability, resulting in a comprehensive evaluation suite of 200 images for the coarse-grained set and 360 images for the fine-grained set.
It is important to note that while previous methods like ColorPeel require separate training procedures for each new target color, our ColorWave approach inherently accepts any arbitrary RGB value without additional training. The distinction between “coarse” and “fine” grained colors is only maintained here for comparative purposes — our method treats all colors equally, handling any point in the RGB color space with the same level of precision. To demonstrate this capability, we include additional experiments with randomly selected colors outside both predefined sets, showing that ColorWave maintains consistent performance across the entire color spectrum without any per-color optimization.
Evaluation metrics. We adopt the evaluation metrics proposed by ColorPeel [4]. Specifically, we compute several complementary metrics:
@@
For metric and implementation details refer to appendix.
Comparison methods. Our method ColorWave represents the first training-free approach for precise color control in diffusion models, establishing a distinct methodological category from existing approaches that require color-specific training. Our primary comparison is with ColorPeel [4], the current state-of-the-art in color-specific prompt learning. For completeness, we compare against all baselines evaluated in ColorPeel, that includes training-free T2I generation baselines: (1) vanilla Stable Diffusion with color name text prompts; (2) Rich-Text [11], which enhances adherence to complex text descriptions; And also training-based personalization methods besides ColorPeel [4]: (3) Textual Inversion [10], learning new pseudo-words in embedding space; (4) DreamBooth [38], fine-tuning the entire diffusion model; and (5) Custom Diffusion [21], optimizing projection matrices in cross-attention layers. The training-free nature of ColorWave creates a fundamental asymmetry in this comparison – while ColorPeel and other training-based approaches require separate optimization processes for each individual color or complex training regimes to handle multiple colors simultaneously, ColorWave inherently accepts any arbitrary RGB triplet without modification. This represents not just an incremental improvement but a paradigm shift in how precise color control can be achieved in generative models, transcending the limitations of discrete color vocabularies and model-specific optimizations.

### v2

@@ Abstract
Corresponding Author (*): Kai Wang (kai.wang@cityu-dg.edu.cn)
Project page: https://hecoding.github.io/colorwave-page
@@ 1 Introduction
Text-to-image (T2I) diffusion models [deepfloyd, Rombach_2022_CVPR_stablediffusion, saharia2022imagen] have revolutionized image generation, offering unprecedented control over composition [liu2022compositionaldiffusion, wang2024compositional, wang2023dynamic], style [sohn2023styledrop, wang2024instantstyle, xing2024csgo, qin2025freelunch], subject matter [textual_inversion, ruiz2022dreambooth], and visual aesthetics [gomez2024art_illusion, geng2024visual_anagrams]. Among these attributes, color is particularly crucial, fundamentally shaping viewer perception, conveying emotional tone, and serving as a critical design specification in professional contexts from brand identity to product development [singh2006impact]. Despite remarkable progress, T2I models exhibit significant limitations in precise color control. When users specify colors through linguistic descriptors like “red” or “blue”, these terms encompass broad ranges of potential shades, making exact color matching challenging—a capability essential for design applications where color fidelity is non-negotiable.
@@
T2I personalization methods such as DreamBooth [ruiz2022dreambooth] or CustomDiffusion [kumari2022customdiffusion] can address this limitation but often fail to correctly disentangle shape from color information [butt2025colorpeel]. ColorPeel [butt2025colorpeel] enables precise color control by learning specific color prompts through additional training. However, these approaches require separate optimization processes for each individual color, resulting in significant computational overhead and limiting practicality. The need for model fine-tuning creates barriers to flexibly specifying arbitrary colors during inference time. Furthermore, learning-based strategies become increasingly impractical in advanced T2I models with multiple text encoders such as SDXL [podell2023sdxl], SD3 [esser2024sd3], and FLUX [flux2024].
@@
Recent advancements in image conditioning for diffusion models, particularly through mechanisms like IP-Adapter [ye2023ip, wang2024instantstyle, xing2024csgo], have enabled new forms of control by conditioning generation on reference images. While these adapters were primarily designed for style transfer or subject-driven generation, they encode complex visual information in ways that have not been fully explored. Through systematic investigation, we discovered a previously unrecognized property: semantic attribute binding which refers to the implicit bindings between visual attributes and their semantic representations within multi-modal space. Specifically, we found that these bindings create a mapping between RGB values in reference images and linguistic color descriptors in text prompts—a connection that existing methods, including IP-Adapter itself, do not explicitly exploit. Our insight enables precise RGB-level color manipulation without any model fine-tuning, revealing capabilities latent within IP-Adapter that haven’t been previously identified or utilized for targeted control.
@@ 2 Related work
Image-Conditioned T2I Diffusion Models. Conditional image-based T2I models integrate auxiliary image inputs with textual prompts, allowing refined control over spatial layout [Li2024ControlNet++:Feedback, Choi2023FineControlNet:Injection], object content [chen2023suti, li2024blipdiffusion], and stylistic attributes [sohn2023styledrop, han2024stylebooth]. ControlNet [zhang2023controlnet] employs a duplicated U-Net [ronneberger2015unet] encoder for pixel-level structural guidance, while UniControlNet [zhao2023unicontrolnet] extends this with dual adapters handling both local and global controls.
@@
As another technique pipeline, IP-Adapter [ye2023ip] leverages the vision encoder from CLIP [radford2021clip] to extract image features, integrating them into the U-Net backbone via cross-attention layers for more coherent conditioning. Building on this, IPAdapter-Instruct [rowles2024ipadapterinstruct] introduces an instructional prompt mechanism, enhancing interpretability and reducing the need for separate model training for different conditioning inputs, addressing a key limitation of ControlNet. However, the role of the additional cross-attention mechanism in IP-Adapter remains underexplored. In this paper, we present the first findings on its properties, revealing that IP-Adapter implicitly establishes a binding between visual attributes and their corresponding semantic representations.
@@
Color Control in T2I diffusion models. With the rapid advancements in T2I generation, various text-guided image editing approaches [hertz2023delta_DDS, meng2022sdedit, mokady2022null, zhang2023forgedit, chen2023fec, li2023stylediffusion, tang2023iterinv] have been developed to enable controllable modifications. For instance, methods like Imagic [kawar2022imagic] and P2P [hertz2022prompt] leverage Stable Diffusion (SD) models for structure-preserving edits. Another technique stream which can also achieve controllable generation is transfer learning for T2I models [ruiz2022dreambooth, kumari2022customdiffusion, wang2024mcti, textual_inversion], or also referred to personalized generation. It aims at adapting a given model to a new concept by given images from the users and bind the new concept with a unique token. As a result, the adaptation model can generate various renditions for the new concept guided by text prompts.
@@
However, all these existing techniques rely heavily on the generative capacity of diffusion models and struggle to achieve fine-grained control over color attributes in image editing and generation tasks. Only a limited number of works [butt2025colorpeel, ge2023richtext] have begun addressing the challenge of precise color generation. Rich-Text [ge2023richtext] enhances color fidelity through multi-pass processing with global and local diffusion models, but at the cost of high computational overhead and reduced color accuracy. In contrast, ColorPeel [butt2025colorpeel] introduces color prompt learning to improve color alignment with user input. However, this method requires extensive training and is restricted to handling only a few color names per training session. In this paper, we argue that the limitations of prior approaches stem from their failure to recognize the implicit attribute binding between the IP-Adapter [ye2023ip] module and T2I diffusion models [podell2023sdxl, Rombach_2022_CVPR_stablediffusion]. Leveraging this property, we propose a training-free solution for precise color generation, allowing seamless adaptation to any user-specified color inputs, particularly benefiting artistic and creative applications. Our work reveals that precise color control capabilities already exist within standard IP-Adapter architectures through semantic attribute binding, eliminating the need for additional training while achieving competitive results.
@@ 3.1 Preliminaries
T2I Diffusion Models. We built on the SDXL [podell2023sdxl] model, consisting of two primary components: an autoencoder and a diffusion model $\epsilon_{\theta}$. The model $\epsilon_{\theta}$ is trained by the loss:
@@
where $\epsilon_{\theta}$ is a UNet, conditioning a latent input $z_{t}$, a timestep $t\sim\text{U}(1,T)$, and a text embedding $\tau_{\xi}(\mathcal{P})$. More specifically, text-guided diffusion models generate an image from the textual condition as $\mathcal{C}_{text}=\tau_{\xi}(\mathcal{P})$, where $\tau_{\xi}$ is the CLIP text encoder [Ilharco_Open_Clip_2021]  [^1] SDXL uses two text encoders and concatenate the embeddings.. The cross-attention map is derived from $\epsilon_{\theta}(z_{t},t,\mathcal{C}_{text})$. After predicting the noise, diffusion schedulers [Lu2022DPM-Solver++:Models, Lu2022DPM-Solver:Steps] are used to predict the latent $z_{t-1}$. As an example with the DDIM scheduler [Song2020DenoisingModels], the formula is:
@@
IP-Adapter. Building on T2I diffusion models, the IP-Adapter [ye2023ip] introduces additional controllability by conditioning the T2I model on a conditional image $\mathcal{I}_{ip}$. Practically, this involves leveraging a pretrained T2I diffusion model and incorporating a cross-attention layer to the (projected) image condition following each text-prompt conditioning layer. The conditional image is encoded in the low-dimensional CLIP image embedding space [Ilharco_Open_Clip_2021] to capture high-level semantic information. By denoting the CLIP image encoder as $\tau_{\phi}$ and IP-Adapter projection as $\mathbf{IP}$, this process is adding a new image condition $\mathcal{C}_{img}=\mathbf{IP}(\tau_{\phi}(\mathcal{I}_{ip}))$ to the T2I model as $z_{t-1}=\mathcal{G}(z_{t},t,\mathcal{C}_{text},\mathcal{C}_{img})$.
@@ 3.2 Semantic attribute binding in IP-Adapters
Figure 2: Illustration of semantic attribute binding. On top, the color-guidance image is provided. a) Given the color name used in the prompt, the generated results will pick the respective color from the color-guidance image. b) Changing the colors of the color guidance image, results in similar changes in the generated images. c) The exact color which should be generated for the used color name can also be a synthetic example.
IP-Adapter [ye2023ip, rowles2024ipadapterinstruct] has demonstrated remarkable effectiveness in transferring visual attributes from reference images to generated content. In this paper, we introduce a key insight: the IP-Adapter implicitly establishes latent-space connections between visual features and their semantic counterparts. As illustrated in Figure 2, this phenomenon, which we call semantic attribute binding, creates a direct correspondence between colors in reference images and their linguistic descriptors in prompts. When using a single reference image containing multiple colors (Figure 2a), the model correctly extracts and applies each color based on the corresponding color name in the prompt (e.g., “green”, “yellow” or “red”). Furthermore, when the colors in the reference image are modified (Figure 2b), the generated results reflect these changes while maintaining the binding to the appropriate color names. This binding even works with synthetic color patches (Figure 2c), allowing for precise RGB-level control. These observations reveal that the adapter inherently associates specific color values with their linguistic descriptors, enabling exact color specification without requiring any additional training or fine-tuning.
Color name attribution. Semantic attribute binding stems from the architectural design of IP-Adapter, particularly its decoupled cross-attention mechanism. As shown by Ye et al. [ye2023ip], the IP-Adapter introduces separate cross-attention layers for image features while keeping the original text cross-attention layers intact. Crucially, both cross-attention mechanisms share the same query matrices, while maintaining separate key and value projection matrices for text and image features:
@@
To effectively leverage this semantic attribute binding, we need to determine which color names in natural language correspond to specific RGB values. Following foundational work in color linguistics by Berlin and Kay [berlin1991basic], we recognize that human languages typically use a limited set of basic color terms to describe the entire color spectrum. In English, there are eleven basic color terms, with eight being chromatic (red, green, blue, yellow, brown, orange, pink, and purple).
@@
Interestingly, we observe that the similarity values are not uniform across all color pairs. Some color terms (like “green” and “blue”) show stronger and more distinctive binding patterns, while others exhibit more diffuse relationships. This variation aligns with findings in color linguistics that certain basic color terms have more centralized and consistent representations across languages and visual systems [kay2006language].
For automatic color name generation, we use a lookup table that maps arbitrary RGB values to optimal linguistic color descriptors. We construct this using traditional color naming models [van2009learning] to select the color name with maximum similarity. This bidirectional mapping between continuous color space and discrete linguistic categories forms the foundation of our color control approach, allowing us to exploit the implicit color knowledge already encoded in the model’s attention mechanisms.
@@
To address these challenges, ColorWave introduces selective attention modulation that preserves generative capabilities while enabling precise color control.
3.3 Spatial Prior Addition
In our approach ColorWave, we leverage the decoupled cross-attention design established in previous adapters [ye2023ip, wang2024instantstyle] for its minimal interference with pretrained knowledge and compatibility with other methods. We implement a strategy of minimal yet precise intervention in the generative model to achieve effective color binding. First, we inject the user-specified color information at a single optimal point in the network, where it most effectively influences the desired subject, following successful approaches in prior work [wang2024instantstyle, wang2024instantstyle_plus]. Second, we rely on the model’s inherent understanding to determine proper color placement based on its pretrained knowledge.
@@ 3.3 Spatial Prior Addition
$\mathbf{x}=\mathbf{A}+\delta_{\lambda}(\tilde{\mathbf{A}}_{o})\mathbf{A}^{\prime}.$ (4)
Here, $\mathbf{A}\in\mathcal{R}^{h\times H\times W}$ represents the original cross-attention output, while $\tilde{\mathbf{A}}_{o}$ is computed by averaging attention maps of the target object across attention heads and $\delta_{\lambda}$ is a thresholding operator preserving the $\lambda$-th percentile. This selective approach complements and ensures effective semantic attribute binding described earlier between the specified color and the target object, while minimizing unintended color attribution to other elements in the scene. The evolution of the object mask throughout the generation process is visualized in Section A.
@@ 4.1 Experimental setup
Dataset. For fair comparison with prior work, we adopt ColorPeel’s [butt2025colorpeel] evaluation framework with two color sets: (1) Coarse-grained: four basic colors (red, green, blue, yellow); (2) Fine-grained: eighteen specialized colors (salmon, beige, navy, indigo, etc.), all with predefined RGB values. We use identical prompts (e.g., “a [color] bowl on the table”) with 20 random seeds each, yielding 200 and 360 images respectively.
Unlike ColorPeel, which requires per-color training, ColorWave accepts arbitrary RGB values without optimization. The coarse/fine distinction is maintained only for comparison—our method treats all colors uniformly across the RGB spectrum. While evaluation focuses on single-object scenarios, ColorWave extends to multi-object scenes (see Section B).
Evaluation metrics. We adopt the evaluation metrics proposed by ColorPeel [butt2025colorpeel]. Specifically, we compute several complementary metrics:
@@
For metric and implementation details refer to Section C.
Comparison methods. ColorWave represents the first training-free approach for precise color control in diffusion models, establishing a distinct methodological category from existing approaches that require color-specific training. Our primary comparison is with ColorPeel [butt2025colorpeel], the current state-of-the-art in color-specific prompt learning. For completeness, we compare against all baselines evaluated in ColorPeel, that includes training-free T2I generation baselines: (1) vanilla Stable Diffusion with color name text prompts; (2) Rich-Text [ge2023richtext] and Rich-Text++ [ge2025expressive], which enhances adherence to complex text descriptions; And also training-based personalization methods besides ColorPeel [butt2025colorpeel]: (3) Textual Inversion [textual_inversion], learning new pseudo-words in embedding space; (4) DreamBooth [ruiz2022dreambooth], fine-tuning the entire diffusion model; and (5) Custom Diffusion [kumari2022customdiffusion], optimizing projection matrices in cross-attention layers. The training-free nature of ColorWave creates a fundamental asymmetry in this comparison – while ColorPeel and other training-based approaches require separate optimization processes for each individual color or complex training regimes to handle multiple colors simultaneously, ColorWave inherently accepts any arbitrary RGB triplet without modification. This represents not just an incremental improvement but a paradigm shift in how to achieve precise color control in generative models, transcending the limitations of discrete color vocabularies and model-specific optimizations.